2019-11-08 ADR Cryostat Run 04 Notes » History » Version 50

« Previous - Version 50/56 (diff) - Next » - Current version
Chris Stoughton, 12/10/2019 04:11 PM

2019-11-08 ADR Cryostat Run 04 Notes

This run was intended as a test of uMUX 5 chip 1 in the dark configuration. We had originally intended to run with uMUX 6 chip 1, but we discovered a short to ground on the flux ramp lines during bonding. It was not clear why this short was present, so in view of the time, we fell back to running with uMUX 5 chip 1. NIST had said this chip was a new design from a batch that seemed to have problems, but we decided to run with it anyway.

2019-11-08 to 2019-11-11 Pump down and leak check

When I left the lab on Friday night, the pressure in the cryostat was 2.31mBar. This was not low enough to start the compressor, so I decided to pump over the weekend. By Monday morning, Chris reported the pressure unchanged at 2.34e-3mBar, but he did not locate a leak when he ran the leak checker.


IN the last few days, we discovered that the magnet current feed through was leaking and HPD sent us a replacement. After fixing that and pumping overnigh, pressure to 2 e-2, turned on PTC and compressor.

Removed He leak checker from system.

Now at 212, 268, 255 deg K with pressure of 1.83e-3

Compressor He pressures are 60 and 290 PSI on the low/high side.


Chris turned on the compressor yesterday, and when I came in today, the cryostat was around 3.2K. The pressure did not seem as good as it should be however: 9.5e-5 mBar. I don't know whether we will be able to cycle the ADR at such a pressure, but I do have some experience with the 3G cryostat that cycling the He10 fridge to 300mK doesn't work very well at pressures around 1e-4 mBar. This is not to say the ADR won't work, but just that I was concerned and wanted to investigate the pressure further.

I turned off the compressor and did a leak check with the cryostat around 10K. There was no obvious leak, but the bottom 300K flange of the cryostat was a little suspicious. It was difficult, however, to distinguish the leak rate here from the secular trend of rising He level due to the increasing pressure of the warming cryostat. After consulting with Chris, I think it makes sense to warm back up to 300K and see what pressure we can achieve at room temperature, possibly coupled with another leak check of the bottom flange of the cryostat. I currently do not understand why we are not able to reach pressures of ~5e-5mBar at room temperature, and much lower at the base temperature of the PTC.

2019-11-18 Leak Check Dewar

9:11 AM PTC is off. Temperatures are 239.0, 193.2, 195.0 pressure 1.18E-3

11:14 AM PTC still off. Temperatures are 241.1, 197.6, 199.3 pressure 1.79E-3

Valve off and look at rate of rise

Up to 3.35e-3 after one minute

1.5 e-2 after 10 minutes.

This smells like a leak!

11:49 Valve off pump. Move hose to leak checker. Pump out hose to 5e-3. Dewar is at 4.2e-3. Now open valve to dewar.

15:12 Dewar at 8.18e-3 He checker at 1.6e-2 at input. Start with partial He pressure of 2.6 e-8. Tent upper area and fill with He. Slow (minutes) rise to 4e-08, open side and clear out He. Rising slower to 6.4e-8. Now up to 7.3e-8 after 10 minutes and sides open. Consistent with slow He leak through viton seals.

15:58 He still rising to 1.7 e-7. Now flood bag over bottom to see if there is a gross leak. He jumped to 3.5 E-6. So, maybe there is a leak around the bottom.
Remove bag and vent.

16:02 Now connect the turbo so we can clear out the He from the seals faster.

16:16 Both pumps connected to dewar via tee. Dewar 5.4e-3 He inlet 8.7 e-3. He partial 2.5e-6 and falling, perhaps due to saturation of viton seals.

16:30 Leave both pumps overnight, at room temperature, to get He out of the seals. Here are the stats:
Partial He: 1.1 e06 with 6.4e-3 at the inlet. Dewar pressure is 1.76e-3. Temperatures: 247.2, 214.7, 217.1

17:24 1.65e-3 mBar, He partial pressure 4.9e-7.

2019-11-18 Check Pumps

Installing the sensitive gauge on the pump, it gets down to 1.29e-5 and falling after 10 minutes. So the pump is pumping.

The other pump complains with an overtemperature warning after a few minutes.


After pumping with He detector and turbo:
8:45 AM
Temperatures: 262.0 238.7 240.3
Partial He 3.3e-9 Inlet 7.1e-3 Dewar 1.59e-3

12:45 PM
Temperatures: 265.0 243.9 245.5
Partial He 2.9e-9 Inlet 7.1e-3 Dewar 1.58e-3

1:00 PM
Temperatures: 265.1 244.1 245.6
Partial He 3.5e-9 Inlet 7.1e-3 Dewar 1.58e-3

He near front side flange shows no response
He near back side flange rose to 4.9e-9 in a minute, then holds steady.

2:00 PM
Bagged lower large flange and front side flange. Will test back side flange. Starting a He partial pressure of 3.2e-9


After pumping overnight:
8:40 AM
Temperatures: 276.3 263.3 264.7
Partial He 1.9e-9 Inlet 7.1e-3 Dewar 1.60e-3

Now throttle valve on main pump to increase he sensitivity. We did not do that yesterday.

Now the Partial He is 3.0e-9 and the dewar is 2.0e-3

Spraying on North side of back-side flange He rises through 1.5e-8 in a minute, levels off at 6e-8. Now open main pump to flush out He.

12:23 PM after pumping with both pumps
Temperatures: 278.1 266.2 267.6
Partial He 4.3e-9 Inlet 7.1e-3 Dewar 1.60e-3

Sealed off the back-side bottom flange. Now try spraying He outside of that seal. Should see no response...unless the leak is somewhere else.
Throttle down the main pump.

12:37 PM with main pump throttled
Temperatures: 278.2 266.4 267.8
Partial He 6.2e-9 Inlet 7.1e-3 Dewar 2.01e-3

After spraying, up to 6.7e-9 after 1 minute, 1.2e-8 after a few minutes. So, this is slower than earlier. The top opening is partially vented, so pump down again, seal the top, and try again.

16:15 Did a spot check of the East side of the bottom flange. Starting He pressure 1.8e-9, No sudden increase at any location. At the end of the test reading 1.1e-8. Reading keep increasing after the test at 16:35 reading is 2e-7 and increasing. Cryostat pressure 1.6e-3.

16:45 He pressure went up to 2.2e-7, and started coming down.


Another day of pumping
3:38 PM
Temperatures: 287.1 281.0 282.3
Partial He 1.3e-9 Inlet 7.9e-3 Dewar 1.57e-3

Close valves; stop He leak check; vent; connect directly to main pump; start PTC

While the dewar was closed off, the pressure rose to 8e-3 in about 5 minutes. That is a rather fast rate of rise.

PTC on. He low/high pressures at 45/295 psi

3:51 PM PTC on
Temperatures: 286.7 281.2 282.3
Dewar 1.77e-3


10:24 AM PTC on overnight
Temperatures: 91.85, 68.46, 62.52
Dewar 1.57e-3
h2. 2019-11-22


1:53 PM PTC on overnight
Temperatures: 66.0, 3.097, 3.092
Dewar 9.43e-5
PTC He pressure LOW/HIGH = 60/260


8:50 AM
50K, Magnet, 3k = 65.6, 3.074, 3.092
50 mK FAA, 1k GGG, MUX Ch. = 2.87, 2.93, 2.93
Dewar 9.25e-5
PTC He pressure LOW/HIGH = 60/250

Turned on magnet power supply; started new log file adr-log-2019-11-25.csv
Start ramp with delay, ramp up, soak, ramp down of 1, 40, 60, 120 minutes

Close pump out valve to dewar; pressure to 9.6e-6


8:50 AM
50K, Magnet, 3k = 0.149, 0.830, 0.830
50 mK FAA, 1k GGG, MUX Ch. = 65.172, 3.088, 3.088
Dewar 9.30e-5
PTC He pressure LOW/HIGH = 60/260

So this held at 150 mK for 49+ hours.

The supply voltage is falling through 1.28 volts, looks like another hour or so for this cycle.

Temperature reading on SRS 160mK.
Started a new ADR cycle with the same configuration parameters. (1, 40, 60, 120)


9:00 AM, found the CryoBoss program not running. Restarted it.
50K, Magnet, 3k = 1.667, 1.1135, 1.1142
50 mK FAA, 1k GGG, MUX Ch. = 63.914, 3.073, 3.063
Dewar 9.60e-5
PTC He pressure LOW/HIGH = 60/250

Start ADR cycle

1:30 -- regulating nicely

50K, Magnet, 3k = 0.15026, 0.615, 0.615
50 mK FAA, 1k GGG, MUX Ch. = 65.469, 3.061, 3.127
Dewar 9.30e-5
PTC He pressure LOW/HIGH = 60/250


11:20 AM

There is a pop up message "Not enough memory to complete this operation" on CryoBoss
Started new ADR cycle

Pressure: 9,2e-5

55K, 2.1K 2.8K


8:53 AM -- temperature at 180 mK, so this cycle lasted about 38 hours. Stop regulating, wait for magnet current and voltage to drop.

50K, Magnet, 3k = 0.184, 0.803, 0.800
50 mK FAA, 1k GGG, MUX Ch. = 55.449, 2.859, 2.871
Dewar 9.50e-5
PTC He pressure LOW/HIGH = 60/250

8:57 AM Start cycle

1:33 PM -- regulating, but electricians need to work on panel, so stop regulating, wait the I and V to get less than 0.05, and turn off power supply
50K, Magnet, 3k = 0.043, 0.622, 0.622
50 mK FAA, 1k GGG, MUX Ch. = 57.175, 2.906, 2.953
Dewar 9.20e-5
PTC He pressure LOW/HIGH = 60/260

4:33 PM -- No electrical work today. Magnet is off, but things are still cold
50K, Magnet, 3k = 0.043, 0.622, 0.622
50 mK FAA, 1k GGG, MUX Ch. = 57.175, 2.901, 2.947
Dewar 9.30e-5
PTC He pressure LOW/HIGH = 60/260


With no regulation over the weekend, temperature up to 0.7 K

50K, Magnet, 3k = -.723, 0.956, 0.956
50 mK FAA, 1k GGG, MUX Ch. = 56.3, 2.899, 2.945
Dewar 8.90e-5
PTC He pressure LOW/HIGH = 60/260

Start an ADR cycle


Still running, nice resonances, trouble finding signal ramp.

4:11 PM

50K, Magnet, 3k = 0.148, 0.740, 0.740
50 mK FAA, 1k GGG, MUX Ch. = 55.4, 2.847, 2.887
Dewar 9.40e-5
PTC He pressure LOW/HIGH = 60/250

Stop temperature control, turn off magnet